# QUEST#4X: an extension of QUEST#4 for benchmarking multireference wavefunction methods

### v1

@@ 4.4 Vertical Excitation Energies
| Radical | State | X-TDA | SS-NEVPT2 | MS-NEVPT2 | SDSPT2 | SDSCI | ic-MRCISD | iCIPT2 |
| allyl | -0.24 | 0.19 | 0.19 | 0.16 | -0.15 | -0.06 | 4.94 | |
| (C3H5) | -0.28 | 0.24 | 0.24 | 0.21 | -0.11 | -0.03 | 5.46 | |
| -0.36 | 0.18 | 0.18 | 0.16 | -0.17 | -0.08 | 5.61 | ||
| 0.07 | 0.16 | 0.15 | 0.10 | -0.13 | -0.05 | 3.40 | ||
| -0.39 | 0.41 | 0.07 | 0.22 | 0.03 | 0.08 | 5.71 | ||
| -0.60 | 0.47 | 0.12 | 0.15 | 0.07 | 0.07 | 6.53 | ||
| -1.52 | -0.64 | 0.07 | -0.07 | -0.08 | -0.11 | 7.80 | ||
| -0.31 | 0.22 | 0.22 | 0.19 | -0.25 | -0.12 | 5.98 | ||
| -0.50 | 0.24 | 0.24 | 0.21 | -0.27 | -0.14 | 7.68 | ||
| BeF | -0.33 | 0.01 | 0.01 | 0.01 | -0.03 | -0.03 | 6.27 | |
| 0.05 | 0.02 | 0.02 | 0.04 | 0.06 | 0.03 | 4.15 | ||
| BeH | -0.37 | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 5.52 | |
| 0.45 | 0.19 | 0.19 | 0.19 | 0.14 | 0.00 | 5.72 | ||
| 0.08 | 0.03 | 0.03 | 0.03 | 0.01 | 0.00 | 2.49 | ||
| -0.76 | 0.00 | 0.00 | 0.01 | 0.01 | 0.00 | 6.46 | ||
| 0.44 | 0.24 | 0.23 | 0.23 | 0.16 | 0.01 | 7.34 | ||
| 0.12 | 0.19 | 0.20 | 0.21 | 0.11 | -0.03 | 7.80 | ||
| 0.24 | 0.03 | 0.03 | 0.03 | -0.01 | 0.00 | 5.88 | ||
| BH2 | -0.24 | -0.03 | -0.02 | -0.02 | -0.01 | 0.01 | 5.72 | |
| -0.35 | -0.04 | -0.03 | -0.03 | -0.02 | 0.01 | 6.75 | ||
| 0.17 | -0.03 | -0.02 | -0.02 | 0.00 | 0.01 | 6.41 | ||
| 0.08 | 0.06 | 0.06 | 0.06 | 0.04 | -0.03 | 6.95 | ||
| 0.19* | 0.05 | 0.05 | 0.06 | 0.04 | 0.00 | 6.38 | ||
| -0.40* | -0.04 | -0.03 | -0.04 | -0.03 | 0.00 | 6.58 | ||
| 0.12 | 0.00 | 0.00 | 0.00 | 0.00 | 0.01 | 1.18 | ||
| -0.37 | -0.06 | -0.05 | -0.05 | -0.02 | 0.01 | 6.87 | ||
| 0.19 | -0.03 | -0.02 | -0.02 | 0.00 | 0.01 | 5.35 | ||
| CH | -1.05 | 0.06 | 0.05 | 0.05 | 0.02 | 0.00 | 3.98 | |
| -0.10 | -0.02 | -0.02 | -0.02 | -0.01 | -0.01 | 6.49 | ||
| 0.47 | 0.03 | 0.03 | 0.03 | 0.01 | -0.01 | 2.91 | ||
| 0.10 | 0.02 | 0.02 | 0.03 | 0.01 | -0.01 | 3.29 | ||
| 0.11 | -0.03 | -0.03 | -0.04 | -0.01 | 0.01 | 0.72 | ||
| -0.18 | 0.03 | 0.03 | 0.02 | 0.00 | 0.00 | 7.66 | ||
| CH3 | -0.23 | -0.02 | -0.01 | -0.02 | -0.02 | 0.01 | 5.86 | |
| 0.04* | 0.17* | 0.07 | 0.09 | 0.11 | 0.04 | 6.96 | ||
| -0.45 * | -0.14* | -0.01 | -0.05 | -0.05 | -0.03 | 7.19 | ||
| -0.26 | -0.06 | -0.03 | -0.03 | -0.01 | 0.00 | 7.65 | ||
| CN | 0.36 | 0.02 | 0.04 | 0.05 | 0.02 | -0.01 | 3.22 | |
| -0.46 | 0.04 | 0.05 | 0.05 | 0.00 | -0.01 | 1.34 | ||
| 0.84 | 0.11 | 0.12 | 0.14 | 0.07 | 0.01 | 7.88 | ||
| -0.30 | 0.04 | 0.05 | 0.05 | 0.02 | 0.02 | 6.03 | ||
| -0.70 | 0.07 | 0.08 | 0.09 | 0.03 | 0.00 | 7.12 | ||
| -0.74 | 0.13 | 0.14 | 0.15 | 0.05 | 0.01 | 7.88 | ||
| CNO | 1.49 | -0.08 | -0.08 | -0.08 | 0.05 | 0.01 | 1.62 | |
| 0.33 | 0.10 | 0.11 | 0.11 | 0.06 | 0.08 | 7.63 | ||
| 0.75 | 0.10 | 0.11 | 0.13 | 0.12 | 0.07 | 7.69 | ||
| -0.43 | -0.19 | -0.18 | -0.16 | -0.18 | -0.18 | 7.82 | ||
| 0.31 | 0.04 | 0.05 | 0.05 | 0.02 | 0.02 | 5.48 | ||
| 0.00 | 0.09 | 0.10 | 0.11 | 0.05 | 0.06 | 6.18 | ||
| 0.08 | 0.07 | 0.08 | 0.09 | 0.04 | 0.06 | 5.70 | ||
| CON | 2.42 | -0.04 | -0.04 | -0.04 | 0.03 | 0.00 | 3.80 | |
| -0.47 | 0.04 | 0.04 | 0.07 | 0.07 | 0.05 | 7.15 | ||
| -0.08 | 0.04 | 0.04 | 0.05 | 0.05 | 0.05 | 3.47 | ||
| -0.20 | -0.03 | -0.03 | -0.01 | 0.01 | 0.02 | 2.74 | ||
| CO+ | 0.83 | 0.05 | 0.05 | 0.07 | 0.01 | -0.01 | 5.80 | |
| 0.10 | 0.03 | 0.04 | 0.05 | -0.01 | 0.00 | 3.27 | ||
| -0.27 | 0.06 | 0.07 | 0.08 | 0.02 | 0.01 | 7.27 | ||
| F2BO | 0.15 | 0.04 | 0.04 | 0.04 | 0.01 | 0.01 | 2.77 | |
| 1.25 | 0.21 | 0.21 | 0.23 | 0.04 | 0.11 | 7.79 | ||
| 1.23 | 0.21 | 0.21 | 0.22 | -0.04 | 0.07 | 6.94 | ||
| 1.37 | 0.26 | 0.26 | 0.27 | -0.02 | 0.08 | 6.53 | ||
| 0.07 | -0.01 | -0.01 | -0.01 | -0.01 | 0.00 | 0.70 | ||
| F2BS | 0.09 | 0.10 | 0.10 | 0.10 | -0.01 | 0.04 | 2.94 | |
| 0.09 | 0.08 | 0.08 | 0.10 | 0.16 | 0.12 | 7.02 | ||
| -0.04 | 0.05 | 0.05 | 0.05 | 0.00 | -0.01 | 0.49 | ||
| -0.03 | -0.04 | -0.04 | -0.03 | 0.08 | 0.05 | 6.39 | ||
| H2BO | 0.02 | 0.06 | 0.06 | 0.07 | 0.01 | 0.00 | 3.51 | |
| 1.16 | 0.06 | 0.07 | 0.09 | 0.07 | 0.04 | 7.25 | ||
| 1.03 | 0.00 | 0.00 | 0.00 | 0.04 | 0.02 | 4.61 | ||
| 1.09 | 0.05 | 0.06 | 0.08 | 0.00 | 0.03 | 7.64 | ||
| -0.21 | 0.04 | 0.04 | 0.05 | -0.02 | 0.00 | 2.17 | ||
| 0.98 | 0.08 | 0.08 | 0.11 | 0.07 | 0.05 | 6.13 | ||
| 0.73 | 0.07 | 0.07 | 0.09 | 0.00 | 0.05 | 6.86 | ||
| HCO | 0.24 | 0.17 | 0.17 | 0.17 | -0.05 | 0.03 | 5.45 | |
| 0.07 | 0.17 | 0.18 | 0.18 | 0.07 | 0.08 | 6.15 | ||
| 0.17 | -0.01 | -0.02 | 0.00 | 0.04 | 0.02 | 2.09 | ||
| -0.18 | 0.11 | 0.11 | 0.11 | -0.02 | 0.01 | 7.07 | ||
| -0.28 | 0.08 | 0.10 | 0.11 | -0.01 | 0.00 | 7.73 | ||
| -0.21 | 0.05 | 0.05 | 0.05 | -0.02 | 0.04 | 6.75 | ||
| -0.21 | 0.04 | 0.05 | 0.05 | 0.00 | 0.05 | 6.37 | ||
| HOC | 0.36 | 0.14 | 0.14 | 0.13 | -0.04 | 0.02 | 3.78 | |
| 0.11 | 0.13 | 0.13 | 0.14 | 0.09 | 0.04 | 5.63 | ||
| 0.01 | -0.01 | -0.01 | 0.00 | 0.02 | 0.00 | 0.92 | ||
| 0.20 | 0.07 | 0.07 | 0.10 | 0.06 | 0.03 | 6.13 | ||
| 0.28 | 0.15 | 0.15 | 0.17 | 0.08 | 0.06 | 7.55 | ||
| 0.24 | 0.18 | 0.18 | 0.17 | -0.02 | 0.05 | 7.73 | ||
| 0.05 | -0.08 | -0.08 | -0.08 | 0.00 | 0.02 | 3.84 | ||
| H2PO | -0.05 | 0.18 | 0.24 | 0.25 | 0.08 | 0.04 | 4.21 | |
| 0.41 | 0.08 | 0.18 | 0.20 | 0.06 | 0.05 | 4.78 | ||
| -0.03 | 0.14 | 0.21 | 0.23 | 0.02 | 0.03 | 5.66 | ||
| -0.01 | 0.11 | 0.17 | 0.16 | -0.04 | 0.01 | 2.81 | ||
| 0.37 | 0.63* | 0.41 | 0.44 | 0.30 | 0.21 | 6.20 | ||
| 0.65 | -0.11* | 0.23 | 0.23 | -0.04 | 0.05 | 6.81 | ||
| 0.06 | 0.18 | 0.24 | 0.25 | 0.02 | 0.10 | 7.28 | ||
| 0.57 | 0.27 | 0.32 | 0.32 | 0.10 | 0.18 | 7.64 | ||
| 0.53 | 0.16 | 0.22 | 0.23 | 0.02 | 0.12 | 6.31 | ||
| H2PS | 0.13 | 0.09 | 0.11 | 0.11 | 0.04 | 0.02 | 2.73 | |
| 0.12 | 0.19 | 0.20 | 0.21 | 0.02 | 0.02 | 4.57 | ||
| -0.09 | 0.08 | 0.09 | 0.09 | -0.02 | 0.02 | 1.14 | ||
| 0.59 | 0.12 | 0.13 | 0.14 | 0.04 | 0.09 | 5.42 | ||
| 0.66 | 0.17 | 0.18 | 0.17 | 0.03 | 0.10 | 5.74 | ||
| 0.68* | 0.23 | 0.19 | 0.22 | 0.07 | 0.07 | 6.02 | ||
| -0.76* | 0.34 | 0.40 | 0.40 | 0.15 | 0.17 | 6.76 | ||
| 1.09* | 0.21 | 0.19 | 0.20 | 0.10 | 0.10 | 5.08 | ||
| -0.34* | 0.46 | 0.49 | 0.51 | 0.23 | 0.22 | 5.73 | ||
| -0.05 | 0.14 | 0.16 | 0.16 | -0.02 | 0.06 | 6.72 | ||
| NCO | 0.22 | -0.04 | -0.03 | -0.03 | 0.05 | 0.01 | 2.87 | |
| 1.22 | 0.07 | 0.08 | 0.08 | 0.03 | 0.03 | 7.03 | ||
| 0.88 | 0.03 | 0.05 | 0.05 | 0.02 | 0.02 | 4.72 | ||
| 0.04 | 0.10 | 0.11 | 0.12 | 0.07 | 0.05 | 7.43 | ||
| -0.12 | 0.06 | 0.07 | 0.08 | 0.04 | 0.06 | 6.79 | ||
| NH2 | 0.08 | 0.02 | 0.02 | 0.02 | 0.00 | 0.00 | 2.12 | |
| 0.04 | 0.06 | 0.06 | 0.05 | 0.00 | 0.00 | 7.69 | ||
| 0.01 | 0.06 | 0.06 | 0.07 | 0.01 | 0.01 | 6.48 | ||
| 0.02 | 0.03 | 0.03 | 0.02 | -0.02 | 0.00 | 7.75 | ||
| 0.04 | 0.04 | 0.04 | 0.03 | 0.00 | 0.01 | 7.29 | ||
| nitromethyl | 0.85 | -0.03 | 0.04 | 0.00 | 0.22 | 0.16 | 2.49 | |
| (CH2NO2) | 0.35 | -0.04 | 0.02 | 0.00 | 0.06 | 0.08 | 5.19 | |
| 1.13 | -0.05 | -0.02 | -0.04 | 0.10 | 0.08 | 2.30 | ||
| 0.63 | 0.02 | 0.12 | 0.14 | 0.10 | 0.14 | 6.71 | ||
| 1.04 | -0.01 | 0.05 | 0.03 | 0.15 | 0.13 | 1.98 | ||
| 0.35 | 0.00 | 0.06 | 0.04 | 0.01 | 0.05 | 4.64 | ||
| 0.23 | -0.04 | 0.09 | 0.14 | 0.44 | 0.32 | 5.30 | ||
| 0.05 | 0.00 | 0.07 | 0.05 | 0.07 | 0.09 | 4.94 | ||
| -0.50 | -0.03 | 0.03 | 0.00 | -0.03 | 0.01 | 4.28 | ||
| 0.10 | 0.01 | 0.08 | 0.05 | 0.01 | 0.06 | 4.45 | ||
| NO | 0.46 | 0.19 | 0.19 | 0.19 | -0.09 | -0.01 | 6.11 | |
| 0.28 | 0.17 | 0.17 | 0.16 | -0.11 | -0.02 | 7.28 | ||
| -0.45 | -0.04 | -0.04 | -0.04 | -0.15 | -0.13 | 7.79 | ||
| -0.21 | -0.04 | -0.04 | -0.02 | 0.05 | 0.02 | 6.36 | ||
| -0.45 | 0.06 | 0.06 | 0.05 | -0.03 | 0.02 | 6.78 | ||
| OH | 0.04 | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 4.11 | |
| 0.09 | 0.01 | 0.01 | 0.01 | -0.01 | -0.01 | 8.03 | ||
| 0.09 | 0.02 | 0.02 | 0.02 | 0.00 | -0.01 | 7.50 | ||
| PH2 | 0.05 | 0.07 | 0.07 | 0.07 | 0.04 | -0.01 | 2.76 | |
| -0.15 | 0.05 | 0.05 | 0.04 | 0.00 | 0.02 | 5.90 | ||
| 0.24 | 0.21 | 0.21 | 0.22 | 0.10 | 0.02 | 7.10 | ||
| 0.08 | 0.10 | 0.10 | 0.10 | 0.03 | -0.02 | 5.17 | ||
| -0.07 | 0.08 | 0.09 | 0.09 | 0.02 | -0.02 | 5.75 | ||
| -0.14 | 0.06 | 0.06 | 0.06 | 0.01 | 0.01 | 7.47 | ||
| 0.09 | 0.04 | 0.04 | 0.04 | 0.00 | 0.03 | 6.16 | ||
| -0.03 | 0.13 | 0.13 | 0.13 | 0.06 | 0.05 | 6.95 | ||
| vinyl | 0.75* | 0.10 | 0.09 | 0.11 | 0.02 | 0.02 | 5.60 | |
| (C2H3) | -0.80* | 0.13 | 0.17 | 0.15 | -0.06 | 0.01 | 6.19 | |
[…]

### v2

@@ 4.4 Vertical Excitation Energies
Radical | State | X-TDA | SS-NEVPT2 | MS-NEVPT2 | SDSPT2 | SDSCI | ic-MRCISD | iCIPT2a | 95% intervalb |
Allyl | -0.24 | 0.19 | 0.19 | 0.16 | -0.15 | -0.06 | 4.94 0.04 | 0.03 | |
(C3H5) | -0.28 | 0.24 | 0.24 | 0.21 | -0.11 | -0.03 | 5.46 0.06 | 0.03 | |
-0.36 | 0.18 | 0.18 | 0.16 | -0.17 | -0.08 | 5.61 0.01 | 0.03 | ||
0.07 | 0.16 | 0.15 | 0.10 | -0.13 | -0.05 | 3.40 0.02 | 0.03 | ||
-0.39 | 0.41 | 0.07 | 0.22 | 0.03 | 0.08 | 5.71 0.05 | 0.02 | ||
-0.60 | 0.47 | 0.12 | 0.15 | 0.07 | 0.07 | 6.53 0.01 | 0.03 | ||
-1.52 | -0.64 | 0.07 | -0.07 | -0.08 | -0.11 | 7.80 0.02 | 0.04 | ||
-0.31 | 0.22 | 0.22 | 0.19 | -0.25 | -0.12 | 5.98 0.02 | 0.02 | ||
-0.50 | 0.24 | 0.24 | 0.21 | -0.27 | -0.14 | 7.68 0.02 | 0.02 | ||
BeFc | -0.33 | 0.01 | 0.01 | 0.01 | -0.03 | -0.03 | 6.27 0.00 | 0.01 | |
0.05 | 0.02 | 0.02 | 0.04 | 0.06 | 0.03 | 4.15 0.00 | 0.00 | ||
BeHc | -0.37 | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 5.52 0.00 | 0.00 | |
0.45 | 0.19 | 0.19 | 0.19 | 0.14 | 0.00 | 5.72 0.00 | 0.00 | ||
0.08 | 0.03 | 0.03 | 0.03 | 0.01 | 0.00 | 2.49 0.00 | 0.00 | ||
-0.76 | 0.00 | 0.00 | 0.01 | 0.01 | 0.00 | 6.46 0.00 | 0.00 | ||
0.44 | 0.24 | 0.23 | 0.23 | 0.16 | 0.01 | 7.34 0.00 | 0.00 | ||
0.12 | 0.19 | 0.20 | 0.21 | 0.11 | -0.03 | 7.80 0.00 | 0.00 | ||
0.24 | 0.03 | 0.03 | 0.03 | -0.01 | 0.00 | 5.88 0.00 | 0.00 | ||
BH2 | -0.24 | -0.03 | -0.02 | -0.02 | -0.01 | 0.01 | 5.72 0.00 | 0.00 | |
-0.35 | -0.04 | -0.03 | -0.03 | -0.02 | 0.01 | 6.75 0.00 | 0.00 | ||
0.17 | -0.03 | -0.02 | -0.02 | 0.00 | 0.01 | 6.41 0.00 | 0.00 | ||
0.08 | 0.06 | 0.06 | 0.06 | 0.04 | -0.03 | 6.95 0.00 | 0.00 | ||
0.19d | 0.05 | 0.05 | 0.06 | 0.04 | 0.00 | 6.38 0.00 | 0.00 | ||
-0.40d | -0.04 | -0.03 | -0.04 | -0.03 | 0.00 | 6.58 0.00 | 0.00 | ||
0.12 | 0.00 | 0.00 | 0.00 | 0.00 | 0.01 | 1.18 0.00 | 0.00 | ||
-0.37 | -0.06 | -0.05 | -0.05 | -0.02 | 0.01 | 6.87 0.00 | 0.00 | ||
0.19 | -0.03 | -0.02 | -0.02 | 0.00 | 0.01 | 5.35 0.00 | 0.00 | ||
CH | -1.05 | 0.06 | 0.05 | 0.05 | 0.02 | 0.00 | 3.98 0.00 | 0.00 | |
-0.10 | -0.02 | -0.02 | -0.02 | -0.01 | -0.01 | 6.49 0.00 | 0.00 | ||
0.47 | 0.03 | 0.03 | 0.03 | 0.01 | -0.01 | 2.91 0.00 | 0.00 | ||
0.10 | 0.02 | 0.02 | 0.03 | 0.01 | -0.01 | 3.29 0.00 | 0.00 | ||
0.11 | -0.03 | -0.03 | -0.04 | -0.01 | 0.01 | 0.72 0.00 | 0.00 | ||
-0.18 | 0.03 | 0.03 | 0.02 | 0.00 | 0.00 | 7.66 0.00 | 0.00 | ||
CH3 | -0.23 | -0.02 | -0.01 | -0.02 | -0.02 | 0.01 | 5.86 0.00 | 0.00 | |
0.04d | 0.17d | 0.07 | 0.09 | 0.11 | 0.04 | 6.96 0.00 | 0.00 | ||
-0.45d | -0.14d | -0.01 | -0.05 | -0.05 | -0.03 | 7.19 0.00 | 0.00 | ||
-0.26 | -0.06 | -0.03 | -0.03 | -0.01 | 0.00 | 7.65 0.00 | 0.00 | ||
CN | 0.36 | 0.02 | 0.04 | 0.05 | 0.02 | -0.01 | 3.22 0.00 | 0.00 | |
-0.46 | 0.04 | 0.05 | 0.05 | 0.00 | -0.01 | 1.34 0.00 | 0.00 | ||
0.84 | 0.11 | 0.12 | 0.14 | 0.07 | 0.01 | 7.88 0.00 | 0.00 | ||
-0.30 | 0.04 | 0.05 | 0.05 | 0.02 | 0.02 | 6.03 0.00 | 0.00 | ||
-0.70 | 0.07 | 0.08 | 0.09 | 0.03 | 0.00 | 7.12 0.00 | 0.00 | ||
-0.74 | 0.13 | 0.14 | 0.15 | 0.05 | 0.01 | 7.88 0.00 | 0.00 | ||
CNO | 1.49 | -0.08 | -0.08 | -0.08 | 0.05 | 0.01 | 1.62 0.01 | 0.00 | |
0.33 | 0.10 | 0.11 | 0.11 | 0.06 | 0.08 | 7.63 0.03 | 0.01 | ||
0.75 | 0.10 | 0.11 | 0.13 | 0.12 | 0.07 | 7.69 0.02 | 0.01 | ||
-0.43 | -0.19 | -0.18 | -0.16 | -0.18 | -0.18 | 7.82 0.02 | 0.01 | ||
0.31 | 0.04 | 0.05 | 0.05 | 0.02 | 0.02 | 5.48 0.01 | 0.00 | ||
0.00 | 0.09 | 0.10 | 0.11 | 0.05 | 0.06 | 6.18 0.03 | 0.01 | ||
0.08 | 0.07 | 0.08 | 0.09 | 0.04 | 0.06 | 5.70 0.01 | 0.00 | ||
CON | 2.42 | -0.04 | -0.04 | -0.04 | 0.03 | 0.00 | 3.80 0.02 | 0.01 | |
-0.47 | 0.04 | 0.04 | 0.07 | 0.07 | 0.05 | 7.15 0.02 | 0.01 | ||
-0.08 | 0.04 | 0.04 | 0.05 | 0.05 | 0.05 | 3.47 0.01 | 0.00 | ||
-0.20 | -0.03 | -0.03 | -0.01 | 0.01 | 0.02 | 2.74 0.00 | 0.00 | ||
CO+ | 0.83 | 0.05 | 0.05 | 0.07 | 0.01 | -0.01 | 5.80 0.00 | 0.00 | |
0.10 | 0.03 | 0.04 | 0.05 | -0.01 | 0.00 | 3.27 0.00 | 0.00 | ||
-0.27 | 0.06 | 0.07 | 0.08 | 0.02 | 0.01 | 7.27 0.00 | 0.00 | ||
F2BO | 0.15 | 0.04 | 0.04 | 0.04 | 0.01 | 0.01 | 2.77 0.01 | 0.04 | |
1.25 | 0.21 | 0.21 | 0.23 | 0.04 | 0.11 | 7.79 0.12 | 0.04 | ||
1.23 | 0.21 | 0.21 | 0.22 | -0.04 | 0.07 | 6.94 0.07 | 0.03 | ||
1.37 | 0.26 | 0.26 | 0.27 | -0.02 | 0.08 | 6.53 0.11 | 0.03 | ||
0.07 | -0.01 | -0.01 | -0.01 | -0.01 | 0.00 | 0.70 0.04 | 0.03 | ||
F2BS | 0.09 | 0.09 | 0.09 | 0.10 | -0.01 | 0.04 | 2.95 0.02 | 0.05 | |
0.09 | 0.08 | 0.08 | 0.10 | 0.16 | 0.12 | 7.02 0.05 | 0.04 | ||
-0.04 | 0.05 | 0.05 | 0.05 | -0.01 | -0.02 | 0.50 0.00 | 0.04 | ||
-0.03 | -0.04 | -0.04 | -0.03 | 0.08 | 0.05 | 6.39 0.03 | 0.05 | ||
H2BO | 0.02 | 0.06 | 0.06 | 0.07 | 0.01 | 0.00 | 3.51 0.00 | 0.01 | |
1.16 | 0.06 | 0.07 | 0.09 | 0.07 | 0.04 | 7.25 0.00 | 0.01 | ||
1.03 | 0.00 | 0.00 | 0.00 | 0.04 | 0.02 | 4.61 0.00 | 0.01 | ||
1.09 | 0.05 | 0.06 | 0.08 | 0.00 | 0.03 | 7.64 0.00 | 0.01 | ||
-0.21 | 0.04 | 0.04 | 0.05 | -0.02 | 0.00 | 2.17 0.00 | 0.01 | ||
0.98 | 0.08 | 0.08 | 0.11 | 0.07 | 0.05 | 6.13 0.01 | 0.01 | ||
0.73 | 0.07 | 0.07 | 0.09 | 0.00 | 0.05 | 6.86 0.00 | 0.01 | ||
HCO | 0.24 | 0.17 | 0.17 | 0.17 | -0.05 | 0.03 | 5.45 0.01 | 0.01 | |
0.07 | 0.17 | 0.18 | 0.18 | 0.07 | 0.08 | 6.15 0.01 | 0.01 | ||
0.17 | -0.01 | -0.02 | 0.00 | 0.04 | 0.02 | 2.09 0.00 | 0.01 | ||
-0.18 | 0.11 | 0.11 | 0.11 | -0.02 | 0.01 | 7.07 0.01 | 0.01 | ||
-0.28 | 0.08 | 0.10 | 0.11 | -0.01 | 0.00 | 7.73 0.01 | 0.01 | ||
-0.21 | 0.05 | 0.05 | 0.05 | -0.02 | 0.04 | 6.75 0.01 | 0.01 | ||
-0.21 | 0.04 | 0.05 | 0.05 | 0.00 | 0.05 | 6.37 0.01 | 0.01 | ||
HOC | 0.36 | 0.14 | 0.14 | 0.13 | -0.04 | 0.02 | 3.78 0.00 | 0.01 | |
0.11 | 0.13 | 0.13 | 0.14 | 0.09 | 0.04 | 5.63 0.01 | 0.01 | ||
0.01 | -0.01 | -0.01 | 0.00 | 0.02 | 0.00 | 0.92 0.00 | 0.01 | ||
0.20 | 0.07 | 0.07 | 0.10 | 0.06 | 0.03 | 6.13 0.01 | 0.01 | ||
0.28 | 0.15 | 0.15 | 0.17 | 0.08 | 0.06 | 7.55 0.01 | 0.01 | ||
0.24 | 0.18 | 0.18 | 0.17 | -0.02 | 0.05 | 7.73 0.01 | 0.01 | ||
0.05 | -0.08 | -0.08 | -0.08 | 0.00 | 0.02 | 3.84 0.01 | 0.01 | ||
H2PO | -0.05 | 0.18 | 0.24 | 0.25 | 0.08 | 0.04 | 4.21 0.01 | 0.03 | |
0.41 | 0.08 | 0.18 | 0.20 | 0.06 | 0.05 | 4.78 0.01 | 0.04 | ||
-0.03 | 0.14 | 0.21 | 0.23 | 0.02 | 0.03 | 5.66 0.02 | 0.02 | ||
-0.01 | 0.11 | 0.17 | 0.16 | -0.04 | 0.01 | 2.81 0.01 | 0.03 | ||
0.37 | 0.63d | 0.41 | 0.44 | 0.30 | 0.21 | 6.20 0.02 | 0.04 | ||
0.65 | -0.11d | 0.23 | 0.23 | -0.04 | 0.05 | 6.81 0.03 | 0.03 | ||
0.06 | 0.18 | 0.24 | 0.25 | 0.02 | 0.10 | 7.28 0.02 | 0.03 | ||
0.57 | 0.27 | 0.32 | 0.32 | 0.10 | 0.18 | 7.64 0.02 | 0.03 | ||
0.53 | 0.16 | 0.22 | 0.23 | 0.02 | 0.12 | 6.31 0.02 | 0.03 | ||
H2PS | 0.13 | 0.09 | 0.11 | 0.11 | 0.04 | 0.02 | 2.73 0.01 | 0.01 | |
0.12 | 0.19 | 0.20 | 0.21 | 0.02 | 0.02 | 4.57 0.02 | 0.01 | ||
-0.09 | 0.08 | 0.09 | 0.09 | -0.02 | 0.02 | 1.14 0.00 | 0.01 | ||
0.59 | 0.12 | 0.13 | 0.14 | 0.04 | 0.09 | 5.42 0.03 | 0.01 | ||
0.66 | 0.17 | 0.18 | 0.17 | 0.03 | 0.10 | 5.74 0.01 | 0.01 | ||
0.68d | 0.23 | 0.19 | 0.22 | 0.07 | 0.07 | 6.02 0.01 | 0.01 | ||
-0.76d | 0.34 | 0.40 | 0.40 | 0.15 | 0.17 | 6.76 0.02 | 0.01 | ||
1.09d | 0.21 | 0.19 | 0.20 | 0.10 | 0.10 | 5.08 0.01 | 0.02 | ||
-0.34d | 0.46 | 0.49 | 0.51 | 0.23 | 0.22 | 5.73 0.02 | 0.01 | ||
-0.05 | 0.14 | 0.16 | 0.16 | -0.02 | 0.06 | 6.72 0.01 | 0.01 | ||
NCO | 0.22 | -0.04 | -0.03 | -0.03 | 0.05 | 0.01 | 2.87 0.01 | 0.03 | |
1.22 | 0.07 | 0.08 | 0.08 | 0.03 | 0.03 | 7.03 0.02 | 0.02 | ||
0.88 | 0.03 | 0.05 | 0.05 | 0.02 | 0.02 | 4.72 0.01 | 0.02 | ||
0.04 | 0.10 | 0.11 | 0.12 | 0.07 | 0.05 | 7.43 0.03 | 0.02 | ||
-0.12 | 0.06 | 0.07 | 0.08 | 0.04 | 0.06 | 6.79 0.01 | 0.02 | ||
NH2 | 0.08 | 0.02 | 0.02 | 0.02 | 0.00 | 0.00 | 2.12 0.00 | 0.00 | |
0.04 | 0.06 | 0.06 | 0.05 | 0.00 | 0.00 | 7.69 0.00 | 0.00 | ||
0.01 | 0.06 | 0.06 | 0.07 | 0.01 | 0.01 | 6.48 0.00 | 0.00 | ||
0.02 | 0.03 | 0.03 | 0.02 | -0.02 | 0.00 | 7.75 0.00 | 0.00 | ||
0.04 | 0.04 | 0.04 | 0.03 | 0.00 | 0.01 | 7.29 0.00 | 0.00 | ||
Nitromethyl | 0.84 | -0.03 | 0.03 | -0.01 | 0.21 | 0.15 | 2.50 0.01 | 0.03 | |
(CH2NO2) | 0.35 | -0.05 | 0.01 | -0.01 | 0.06 | 0.07 | 5.20 0.01 | 0.04 | |
1.13 | -0.05 | -0.02 | -0.04 | 0.10 | 0.08 | 2.30 0.00 | 0.03 | ||
0.64 | 0.03 | 0.13 | 0.15 | 0.11 | 0.15 | 6.70 0.07 | 0.03 | ||
1.02 | -0.03 | 0.03 | 0.01 | 0.13 | 0.11 | 2.00 0.02 | 0.04 | ||
0.33 | -0.02 | 0.04 | 0.02 | -0.01 | 0.03 | 4.66 0.02 | 0.04 | ||
0.21 | -0.05 | 0.08 | 0.13 | 0.42 | 0.30 | 5.31 0.00 | 0.03 | ||
0.04 | 0.00 | 0.06 | 0.04 | 0.06 | 0.09 | 4.94 0.03 | 0.02 | ||
-0.52 | -0.05 | 0.01 | -0.02 | -0.05 | 0.00 | 4.30 0.02 | 0.03 | ||
0.10 | 0.01 | 0.07 | 0.05 | 0.00 | 0.05 | 4.46 0.03 | 0.04 | ||
NO | 0.46 | 0.19 | 0.19 | 0.19 | -0.09 | -0.01 | 6.11 0.01 | 0.01 | |
0.28 | 0.17 | 0.17 | 0.16 | -0.11 | -0.02 | 7.28 0.01 | 0.00 | ||
-0.45 | -0.04 | -0.04 | -0.04 | -0.15 | -0.13 | 7.79 0.00 | 0.01 | ||
-0.21 | -0.04 | -0.04 | -0.02 | 0.05 | 0.02 | 6.36 0.00 | 0.01 | ||
-0.45 | 0.06 | 0.06 | 0.05 | -0.03 | 0.02 | 6.78 0.00 | 0.01 | ||
OH | 0.04 | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 4.11 0.00 | 0.00 | |
0.09 | 0.01 | 0.01 | 0.01 | -0.01 | -0.01 | 8.03 0.00 | 0.00 | ||
0.09 | 0.02 | 0.02 | 0.02 | 0.00 | -0.01 | 7.50 0.00 | 0.00 | ||
PH2 | 0.05 | 0.07 | 0.07 | 0.07 | 0.04 | -0.01 | 2.76 0.00 | 0.00 | |
-0.15 | 0.05 | 0.05 | 0.04 | 0.00 | 0.02 | 5.90 0.00 | 0.00 | ||
0.24 | 0.21 | 0.21 | 0.22 | 0.10 | 0.02 | 7.10 0.00 | 0.00 | ||
0.08 | 0.10 | 0.10 | 0.10 | 0.03 | -0.02 | 5.17 0.00 | 0.00 | ||
-0.07 | 0.08 | 0.09 | 0.09 | 0.02 | -0.02 | 5.75 0.00 | 0.00 | ||
-0.14 | 0.06 | 0.06 | 0.06 | 0.01 | 0.01 | 7.47 0.00 | 0.00 | ||
0.09 | 0.04 | 0.04 | 0.04 | 0.00 | 0.03 | 6.16 0.00 | 0.00 | ||
-0.03 | 0.13 | 0.13 | 0.13 | 0.06 | 0.05 | 6.95 0.00 | 0.00 | ||
Vinyl | 0.75d | 0.10 | 0.09 | 0.11 | 0.02 | 0.02 | 5.60 0.00 | 0.02 | |
(C2H3) | -0.80d | 0.13 | 0.17 | 0.15 | -0.06 | 0.01 | 6.19 0.01 | 0.01 | |
-0.10 | 0.04 | 0.05 | 0.05 | 0.01 | -0.03 | 3.28 0.00 | 0.01 | ||
[…]